# Reinforcement Learning-Based Market Making as a Stochastic Control on Non-Stationary Limit Order Book Dynamics

### v1

@@ 1 Introduction
Market making in financial markets can be defined as continuously quoting bid and ask prices to profit from thespread, and is an essential part of the market microstructure, as it helps to narrow bid-ask spreads, reduce volatility, and maintain market stability Glosten1985 ; OHara1995 . This is particularly important in times of high uncertainty, as market makers provide liquidity, which helps to maintain a fair and efficient market price. With the advent of electronic trading, placing optimal bid and ask quotes as a market maker is becoming an almost completely automatized task, but such a transition requires dealing with additional caveats, including slippage, market impact, adverse market regimes Cont2010 ; Bouchaud2018 and non-stationary conditions Gasperov2021 .
@@
The Reinforcement Learning (RL) paradigm has shown promising results for optimizing market making strategies, where agents learn to adapt their policies through trial and error given a numerical outcome reward score [^1] Not to be confused with financial returns., which is used to evaluate the agent’s performance in the environment Sutton2018 . The RL approach is based on the Bellman equation for state values or state-action pair values, which recursively define the state-value of a policy as the expected return of discounted future rewards. State-of-the-art RL algorithms, such as Proximal Policy Optimization (PPO) and Soft Actor-Critic (SAC), solve the Bellman equation by approximating the optimal policy using neural networks to learn either the policy, the value function Sutton2018 , or both Schulman2015 ; Mnih2015 , thus enabling the agent to learn the optimal actions according to observed market states He2023 ; Bakshaev2020 .
@@
Using historical data to train RL agents is a common practice in the market making literature, but has some limitations due to being computationally expensive and requiring large amounts of data, besides not including the effects of market impact and inventory risk Frey2023 ; Ganesh2019 on the agent’s decision-making policies. An additional approach to creating realistic environments are agent-based simulations, where generative agents are first trained against observed market messages and then used to simulate order arrivals Jerome2022 ; Selser2021 . This approach has the disadvantage of trading off fine-grained control over market dynamics for realism, as well as limiting the agent’s adaptability to unseen market regimes, which can lead to underfitting Jerome2022 ; Selser2021 and suboptimal decision-making under scenarios where market impact and slippage have a more prominent effect on the agent’s performance. On the other hand, using stochastic models to simulate the limit order book environment is computationally cheaper and faster to train, but may not capture the full complexity of real market dynamics, due to simplified assumptions and static  Sun2022 .
@@
In the context of this paper, we propose and discuss a methodology for integrating a reinforcement learning agent in a market-making context, where the underlying market dynamics are explicitly simulated using parameterizable stochastic processes combined carefully to capture observed stylized facts of real markets. Our main contribution is implementing multiple non-stationary dynamics into a single limit order book simulator, and how using fine-controlled non-stationary environments can enhance the resulting agent’s performance under adverse market conditions and provide a more realistic training environment for RL agents in market-making scenarios, while still considering market impact and inventory risk. We then perform a comparative analysis of the agent’s performance under changing market conditions during the trading day and benchmark it against the Avellaneda-Stoikov market making strategy Avellaneda2008 optimal solution (under a simplified market model).
@@ 2.1 Data Type
The majority of the references used historical data, while the second most common type was agent-based simulations, where generative agents are trained against observed market messages and used to simulate the LOB environment Frey2023 ; Ganesh2019 trading off control over the market dynamics for replicating observed market flow.
@@
The third most common were stationary processes including Geometric Brownian Motion and Poisson processes Gasperov2021 ; Sun2022 for order arrival and price dynamics, while some used the Hawkes process for simulating order arrival times Jerome2022 ; Selser2021 . Overall, the references shown in Figure 1 demonstrate a preference for historical data and model-based simulators, and although being extremely promising approaches, neither completely account for changing market regimes and scenarios where impact and slippage have a more prominent effect on the agent’s performance. With no references using both simulated data and reinforcement learning agents instead of closed-form solutions, we aim to address this gap in the literature by implementing a reinforcement learning agent and testing its robustness and generalization capabilities under non-stationary environments.
@@ 2.2 State, Action, and Reward Spaces
Most analyzed references defined state spaces primarily based on market-level observations, as shown in 2(c). specifically top-of-book quotes and N-depth book levels which align well with the real-time data available to market-making agents He2023 ; Bakshaev2020 . Additionally, agent inventory was also a common feature, reflecting the importance of managing risk and liquidity in market-making strategies Patel2018 ; Ganesh2019 . The tagged action spaces was mostly made up only one pair of quotes, with some references allowing the choice of multiple bid-ask levels. Some references also used book levels as discrete action space variables, displayed in 2(b).
@@ 2.3 Chosen Algorithms
Most references used Deep Q-Learning and Actor-Critic methods, leveraging neural networks for the optimal policy or state-value function estimation, due to the high-dimensional nature of limit order books Patel2018 ; Ganesh2019 . The most used algorithms were PPO and DQN, as shown in Table 1, both being the best performers in terms of convergence and generalization capabilities in the literature Sun2022 ; Gasperov2021a , with some references using DDQN variants. The literature on model-free approaches overall demonstrated strong adaptability capabilities towards changing market conditions while still maintaining acceptable training times. Overall, the performed bibliographical review underscores the growing relevance and preference for reinforcement learning algorithms in the market-making domain, particularly when combined with realistic spatial and temporal state representations.
@@ 3.1.1 Chosen State Space
We choose a state space that tries to best incorporate the historical events of the limit order book into a single observable state using market features deemed relevant and chosen through our initial bibliography research, as well as guided by the works of Gasparov et al. (2021) Gasperov2021 and Gueant et al. (2017) Gueant2017 on reinforcement learning for market-making. We used the agent’s current inventory and a set of indicators for the currently observed market state: the Relative Strength Index (RSI); order imbalance (OI); and micro price (MP). Additionally, for a fixed number $D$ of LOB price levels the pair $(\delta^{d},Q^{d})$, where $\delta^{d}$ is the half-spread distance for the level $d\leq D$, and $Q^{d}$ the amount of orders posted at that level is added to the state as a set of tuples, for both the ask and bid sides of the book. The state space can therefore be represented by the following expression:
@@
where $P_{t}^{\text{ask}}$ and $P_{t}^{\text{bid}}$ represent the best ask and bid prices at time $t$, and $Q_{t}^{\text{bid}}$ and $Q_{t}^{\text{ask}}$ represent the total bid and ask quantities, respectively. The Average Gain and Average Loss are computed over a rolling window (in our case, fixed 5 minute observation intervals), and are defined as the average price increases and decreases during that window, respectively. $\text{OI}_{t}\in[-1,1]$, with $\text{OI}_{t}=1$ indicating complete dominance of bid orders, and $\text{OI}_{t}=-1$ indicating ask order dominance. The RSI is bounded between 0 and 100, with values above 70 indicating overbought conditions and values below 30 indicating oversold conditions. Overall, the chosen state space designed to capture the most relevant market features for the agent to make informed decisions is not a novel approach per-se, containing common indicators used in market-making literature Gueant2022 ; Selser2021a ; FalcesMarin2022 .
@@ 3.1.3 Episodic Reward Function and Returns
The episode reward function $R_{t}\in\mathbb{R}$ reflects the agent’s profit and inventory risk obtained during a specific time in the episode. It depends on the spread and executed quantities, as well as the inventory cost and was choosen according to the best performing reward structures taken from the literature review. For our model the utility chosen is based on a risk-aversion enforced utility function, specifically the constant absolute risk aversion (CARA) Arrow1965 ; Pratt1964 with the virtual running Profit and Loss (PnL) as input. It depends on the realized spread $\delta$ and the realized quantity qq [^2] Differs from the agent’s posted order quantity $Q$, as $q$ is a stochastic variable dependent on the underlying market dynamics., and is computed as follows, where the penalty for holding large inventory positions is discounted by a factor of $\eta$ from the final score:
@@
Finally, the reward function is defined as the negative of the exponential of the running PnL, a common choice for risk-averse utility functions according to literature Gueant2022 ; Selser2021a ; FalcesMarin2022 . The CARA utility function is defined as follows, where $\gamma$ is the risk aversion parameter:
@@ 3.2 State Transition Distribution
In continuous-time and state MDPs, the state dynamics is reflected by $\mathcal{L}$, and modern approaches to optimal control solve analytically by obtaining a closed-form expression for the model’s evolution equations, as originally tackled by Avellaneda et al. (2008) Avellaneda2008 and Gueant et al. (2017) Gueant2017 . or numerically by approximating its transition probabilities Gueant2022 ; Selser2021a ; FalcesMarin2022 . Closed-form expressions for $\mathcal{L}$ are obtainable for simple environment models, by usually disconsidering market impact, which is not the case for our proposed model, and solving for the generator operator is therefore outside the scope of this paper. We describe in section 4 our chosen approach to model the state transition distribution, using the Proximal Policy Optimization (PPO) algorithm to approximate the optimal policy for the agent.
@@ 3.3 Market Model Description and Environment Dynamics
where $\mu$ is the price drift, $s_{t}$ is the mean spread, and $\sigma$ the price volatility. The drift rate process follows a mean-reverting Ornstein-Uhlenbeck process Ornstein1930 , while the spread rate similarly follows a Cox-Ingersoll-Ross process Cox1985 , which, given that the Feller condition $2\kappa_{s}\sigma_{s}^{2}>1$ is satisfied, ensures the spread remains positive:
@@
and while there are no orders on both sides of the book, the midprice is either the last traded price, observed or an initial value set for the simulation, in the given order of priority. By analyzing the midprice process, we can make sure the returns are normally distributed, with mean $\mu_{t}$ and volatility $\frac{\sigma^{2}}{2}$, assuring the model reflects a stylized fact of LOBs commonly observed in markets Gueant2022 .
@@ 3.4 Decision Process
The Bellman equation is used to recursively define state values in terms of immediate rewards and expected value of subsequent states, and for simpler environments, the equation can be solved directly by dynamic programming methods. Such methods, like value iteration or policy iteration, although effective, become computationally expensive and sometimes unfeasible for large state spaces, or require a model of the environment, which is not always available in practice. Neural networks have been successfully used to approximate both the value function or the optimal actions directly, demonstrating state-of-the-art results in various domains, including games, robotics, and finance  He2023 ; Bakshaev2020 ; Patel2018 ; Ganesh2019 ; Sun2022 ; Gasperov2021a . Our simulation environment is complex enough that the state space is too large to store all possible state-action pairs, and since we assume no knowledge of the underlying dynamics of the environment when training the agent, a model-free approach is necessary.
@@ 3.4.1 Generalized Policy Iteration and Policy Gradient
Each pass updates both the policy parameters $\theta$ and the value function parameters $\phi$ by applying the usual backward propagation pass for neural networks to minimize the loss function according to a chosen optimizer. A more detailed explanation of the PPO algorithm can be found in the original paper by Schulman et al. (2017) Schulman2017 . Our implementation of the algorithm is a distributed-parallel version of the original version, where multiple learners interact with the environment in parallel and share the experience to update the policy and value function. Short of the V-trace algorithm, this approach is similar to the IMPALA algorithm Espeholt2018 . Besides the proposed approach for simulating dynamic limit order book environments, we also implement a hybrid neural network for the agent’s policy and value functions, composed of a sequence of self-attention layers to capture the spatial dependencies between the different levels of the LOB, concatenated with dense layers to process the market features, as discussed in section 4.
@@ 3.4.2 Benchmark Closed-Form Expression for Simplified Model
To ensure the agent’s performance is able to capture the complexity of the environment, we use a benchmark closed-form expression to compare the agent’s performance against. We chose the closed-form expression for the optimal bid-ask spread pair as proposed by Avellaneda et al. (2008) Avellaneda2008 , which is given by the following expression:
@@ 4 Implementation and Model Description
To train our policy and value networks, we use the aforementioned PPO algorithmSchulman2017 , which is a model-free, on-policy approach to optimize the policy directly, as discussed in Subsection 3.4.1. Our training loop shown in Algorithm 2 consists of collecting trajectories, computing the Generalized Advantage Estimation (GAE), which is used instead of the usual returns $G_{t}$. The chosen approach can be classified as online reinforcement learning, where the agent learns from interactions with the environment, without the need for a pre-existing dataset, as opposed to offline reinforcement learning, where the agent learns from a static dataset. Additionally, we use a replay buffer to store the trajectories and sample mini-batches for training the policy and value networks. Since the replay buffer is discarded immediately after a small fixed number of epochs, the approach is closer to on-policy learning, as the data collected by the policy is the same as the data used for gradient updates.

### v2

@@ 1 Introduction
Market making in financial markets can be defined as continuously quoting bid and ask prices to profit from thespread, and is an essential part of the market microstructure, as it helps to narrow bid-ask spreads, reduce volatility, and maintain market stability Glosten and Milgrom (1985); O’Hara (1995). This is particularly important in times of high uncertainty, as market makers provide liquidity, which helps to maintain a fair and efficient market price. With the advent of electronic trading, placing optimal bid and ask quotes as a market maker is becoming an almost completely automatized task, but such a transition requires dealing with additional caveats, including slippage, market impact, adverse market regimes Cont et al. (2010); Bouchaud et al. (2018) and non-stationary conditions Gasperov et al. (2021).
@@
The Reinforcement Learning (RL) paradigm has shown promising results for optimizing market making strategies, where agents learn to adapt their policies through trial and error given a numerical outcome reward score [^1] Not to be confused with financial returns., which is used to evaluate the agent’s performance in the environment Sutton and Barto (2018). The RL approach is based on the Bellman equation for state values or state-action pair values, which recursively define the state-value of a policy as the expected return of discounted future rewards. State-of-the-art RL algorithms, such as Proximal Policy Optimization (PPO) and Soft Actor-Critic (SAC), solve the Bellman equation by approximating the optimal policy using neural networks to learn either the policy, the value function Sutton and Barto (2018), or both Schulman et al. (2015); Mnih et al. (2015), thus enabling the agent to learn the optimal actions according to observed market states He et al. (2023); Bakshaev (2020).
@@
Using historical data to train RL agents is a common practice in the market making literature, but has some limitations due to being computationally expensive and requiring large amounts of data, besides not including the effects of market impact and inventory risk Frey et al. (2023); Ganesh et al. (2019) on the agent’s decision-making policies. An additional approach to creating realistic environments are agent-based simulations, where generative agents are first trained against observed market messages and then used to simulate order arrivals. This approach has the disadvantage of trading off fine-grained control over market dynamics for realism, as well as limiting the agent’s adaptability to unseen market regimes, which can lead to underfitting and suboptimal decision-making under scenarios where market impact and slippage have a more prominent effect on the agent’s performance. On the other hand, using stochastic models to simulate the limit order book environment is computationally cheaper and faster to train, but may not capture the full complexity of real market dynamics, due to simplified assumptions and static  Sun et al. (2022).
@@
In the context of this paper, we propose and discuss a methodology for integrating a reinforcement learning agent in a market-making context, where the underlying market dynamics are explicitly simulated using parameterizable stochastic processes combined carefully to capture observed stylized facts of real markets. Our main contribution is implementing multiple non-stationary dynamics into a single limit order book simulator, and how using fine-controlled non-stationary environments can enhance the resulting agent’s performance under adverse market conditions and provide a more realistic training environment for RL agents in market-making scenarios, while still considering market impact and inventory risk. We then perform a comparative analysis of the agent’s performance under changing market conditions during the trading day and benchmark it against the Avellaneda-Stoikov market making strategy Avellaneda and Stoikov (2008) optimal solution (under a simplified market model).
@@ 2.1 Data Type
The majority of the references used historical data, while the second most common type was agent-based simulations, where generative agents are trained against observed market messages and used to simulate the LOB environment Frey et al. (2023); Ganesh et al. (2019) trading off control over the market dynamics for replicating observed market flow.
@@
The third most common were stationary processes including Geometric Brownian Motion and Poisson processes Gasperov et al. (2021); Sun et al. (2022) for order arrival and price dynamics, while some used the Hawkes process for simulating order arrival times. Overall, the references shown in Figure 1 demonstrate a preference for historical data and model-based simulators, and although being extremely promising approaches, neither completely account for changing market regimes and scenarios where impact and slippage have a more prominent effect on the agent’s performance. With no references using both simulated data and reinforcement learning agents instead of closed-form solutions, we aim to address this gap in the literature by implementing a reinforcement learning agent and testing its robustness and generalization capabilities under non-stationary environments.
@@ 2.2 State, Action, and Reward Spaces
Most analyzed references defined state spaces primarily based on market-level observations, as shown in 2(c). specifically top-of-book quotes and N-depth book levels which align well with the real-time data available to market-making agents He et al. (2023); Bakshaev (2020). Additionally, agent inventory was also a common feature, reflecting the importance of managing risk and liquidity in market-making strategies Patel (2018); Ganesh et al. (2019). The tagged action spaces was mostly made up only one pair of quotes, with some references allowing the choice of multiple bid-ask levels. Some references also used book levels as discrete action space variables, displayed in 2(b).
@@ 2.3 Chosen Algorithms
Most references used Deep Q-Learning and Actor-Critic methods, leveraging neural networks for the optimal policy or state-value function estimation, due to the high-dimensional nature of limit order books Patel (2018); Ganesh et al. (2019). The most used algorithms were PPO and DQN, as shown in Table 1, both being the best performers in terms of convergence and generalization capabilities in the literature Sun et al. (2022); Gasperov and Kostanjcar (2021), with some references using DDQN variants. The literature on model-free approaches overall demonstrated strong adaptability capabilities towards changing market conditions while still maintaining acceptable training times. Overall, the performed bibliographical review underscores the growing relevance and preference for reinforcement learning algorithms in the market-making domain, particularly when combined with realistic spatial and temporal state representations.
@@ 3.1.1 Chosen State Space
We choose a state space that tries to best incorporate the historical events of the limit order book into a single observable state using market features deemed relevant and chosen through our initial bibliography research, as well as guided by the works of Gasparov et al. (2021) Gasperov et al. (2021) and Gueant et al. (2017) Guéant (2017) on reinforcement learning for market-making. We used the agent’s current inventory and a set of indicators for the currently observed market state: the Relative Strength Index (RSI); order imbalance (OI); and micro price (MP). Additionally, for a fixed number $D$ of LOB price levels the pair $(\delta^{d},Q^{d})$, where $\delta^{d}$ is the half-spread distance for the level $d\leq D$, and $Q^{d}$ the amount of orders posted at that level is added to the state as a set of tuples, for both the ask and bid sides of the book. The state space can therefore be represented by the following expression:
@@
where $P_{t}^{\text{ask}}$ and $P_{t}^{\text{bid}}$ represent the best ask and bid prices at time $t$, and $Q_{t}^{\text{bid}}$ and $Q_{t}^{\text{ask}}$ represent the total bid and ask quantities, respectively. The Average Gain and Average Loss are computed over a rolling window (in our case, fixed 5 minute observation intervals), and are defined as the average price increases and decreases during that window, respectively. $\text{OI}_{t}\in[-1,1]$, with $\text{OI}_{t}=1$ indicating complete dominance of bid orders, and $\text{OI}_{t}=-1$ indicating ask order dominance. The RSI is bounded between 0 and 100, with values above 70 indicating overbought conditions and values below 30 indicating oversold conditions. Overall, the chosen state space designed to capture the most relevant market features for the agent to make informed decisions is not a novel approach per-se, containing common indicators used in market-making literature Gueant (2022); Selser et al. (2021); Falces Marin et al. (2022).
@@ 3.1.3 Episodic Reward Function and Returns
The episode reward function $R_{t}\in\mathbb{R}$ reflects the agent’s profit and inventory risk obtained during a specific time in the episode. It depends on the spread and executed quantities, as well as the inventory cost and was choosen according to the best performing reward structures taken from the literature review. For our model the utility chosen is based on a risk-aversion enforced utility function, specifically the constant absolute risk aversion (CARA) Arrow (1965); Pratt (1964) with the virtual running Profit and Loss (PnL) as input. It depends on the realized spread $\delta$ and the realized quantity qq [^2] Differs from the agent’s posted order quantity $Q$, as $q$ is a stochastic variable dependent on the underlying market dynamics., and is computed as follows, where the penalty for holding large inventory positions is discounted by a factor of $\eta$ from the final score:
@@
Finally, the reward function is defined as the negative of the exponential of the running PnL, a common choice for risk-averse utility functions according to literature Gueant (2022); Selser et al. (2021); Falces Marin et al. (2022). The CARA utility function is defined as follows, where $\gamma$ is the risk aversion parameter:
@@ 3.2 State Transition Distribution
In continuous-time and state MDPs, the state dynamics is reflected by $\mathcal{L}$, and modern approaches to optimal control solve analytically by obtaining a closed-form expression for the model’s evolution equations, as originally tackled by Avellaneda et al. (2008) Avellaneda and Stoikov (2008) and Gueant et al. (2017) Guéant (2017). or numerically by approximating its transition probabilities Gueant (2022); Selser et al. (2021); Falces Marin et al. (2022). Closed-form expressions for $\mathcal{L}$ are obtainable for simple environment models, by usually disconsidering market impact, which is not the case for our proposed model, and solving for the generator operator is therefore outside the scope of this paper. We describe in section 4 our chosen approach to model the state transition distribution, using the Proximal Policy Optimization (PPO) algorithm to approximate the optimal policy for the agent.
@@ 3.3 Market Model Description and Environment Dynamics
where $\mu$ is the price drift, $s_{t}$ is the mean spread, and $\sigma$ the price volatility. The drift rate process follows a mean-reverting Ornstein-Uhlenbeck process Ornstein and Uhlenbeck (1930), while the spread rate similarly follows a Cox-Ingersoll-Ross process Cox et al. (1985), which, given that the Feller condition $2\kappa_{s}\sigma_{s}^{2}>1$ is satisfied, ensures the spread remains positive:
@@
and while there are no orders on both sides of the book, the midprice is either the last traded price, observed or an initial value set for the simulation, in the given order of priority. By analyzing the midprice process, we can make sure the returns are normally distributed, with mean $\mu_{t}$ and volatility $\frac{\sigma^{2}}{2}$, assuring the model reflects a stylized fact of LOBs commonly observed in markets Gueant (2022).
@@ 3.4 Decision Process
The Bellman equation is used to recursively define state values in terms of immediate rewards and expected value of subsequent states, and for simpler environments, the equation can be solved directly by dynamic programming methods. Such methods, like value iteration or policy iteration, although effective, become computationally expensive and sometimes unfeasible for large state spaces, or require a model of the environment, which is not always available in practice. Neural networks have been successfully used to approximate both the value function or the optimal actions directly, demonstrating state-of-the-art results in various domains, including games, robotics, and finance  He et al. (2023); Bakshaev (2020); Patel (2018); Ganesh et al. (2019); Sun et al. (2022); Gasperov and Kostanjcar (2021). Our simulation environment is complex enough that the state space is too large to store all possible state-action pairs, and since we assume no knowledge of the underlying dynamics of the environment when training the agent, a model-free approach is necessary.
@@ 3.4.1 Generalized Policy Iteration and Policy Gradient
Each pass updates both the policy parameters $\theta$ and the value function parameters $\phi$ by applying the usual backward propagation pass for neural networks to minimize the loss function according to a chosen optimizer. A more detailed explanation of the PPO algorithm can be found in the original paper by Schulman et al. (2017) Schulman et al. (2017). Our implementation of the algorithm is a distributed-parallel version of the original version, where multiple learners interact with the environment in parallel and share the experience to update the policy and value function. Short of the V-trace algorithm, this approach is similar to the IMPALA algorithm Espeholt et al. (2018). Besides the proposed approach for simulating dynamic limit order book environments, we also implement a hybrid neural network for the agent’s policy and value functions, composed of a sequence of self-attention layers to capture the spatial dependencies between the different levels of the LOB, concatenated with dense layers to process the market features, as discussed in section 4.
@@ 3.4.2 Benchmark Closed-Form Expression for Simplified Model
To ensure the agent’s performance is able to capture the complexity of the environment, we use a benchmark closed-form expression to compare the agent’s performance against. We chose the closed-form expression for the optimal bid-ask spread pair as proposed by Avellaneda et al. (2008) Avellaneda and Stoikov (2008), which is given by the following expression:
@@ 4 Implementation and Model Description
To train our policy and value networks, we use the aforementioned PPO algorithmSchulman et al. (2017), which is a model-free, on-policy approach to optimize the policy directly, as discussed in Subsection 3.4.1. Our training loop shown in Algorithm 2 consists of collecting trajectories, computing the Generalized Advantage Estimation (GAE), which is used instead of the usual returns $G_{t}$. The chosen approach can be classified as online reinforcement learning, where the agent learns from interactions with the environment, without the need for a pre-existing dataset, as opposed to offline reinforcement learning, where the agent learns from a static dataset. Additionally, we use a replay buffer to store the trajectories and sample mini-batches for training the policy and value networks. Since the replay buffer is discarded immediately after a small fixed number of epochs, the approach is closer to on-policy learning, as the data collected by the policy is the same as the data used for gradient updates.